# From Code to Correctness: Closing the Last Mile of Code Generation with Hierarchical Debugging

### v2

@@ 4.7 Case Study
We perform a qualitative analysis of how MGDebugger effectively identifies and corrects buggy parts compared to baseline methods. Figure 5 shows an example of debugging code snippets from the HumanEvalFix dataset using MGDebugger and other representative methods, with DeepSeek-Coder-V2-Lite as the backbone LLM. The original buggy solution computes the Collatz sequence with an incorrect computation logic of $n=n\times 2+1$. While other methods correct the computation to $n=n\times 3+1$, they introduce a new bug that misses the last “1” in the Collatz sequence. This is possibly because they get distracted by the need to filter odd numbers, and thus move the operation of appending the number to the results before updating $n$. MGDebugger excelled by decomposing the problem into distinct subfunctions: sequence generation and odd number filtering. By debugging each subfunction independently, MGDebugger ensured comprehensive error correction, including the subtle requirement of incorporating 1 into the Collatz sequence. This approach demonstrates MGDebugger’s ability to handle complex, multi-step problems more effectively than holistic debugging methods. Additionally, it highlights MGDebugger’s ability to not only fix bugs but also restructure code for enhanced clarity and correctness, demonstrating its potential in improving the quality of LLM-generated code. More examples and analysis on the three datasets can be found in Appendix F.
5 Conclusion
@@
Future work can build upon this foundation to develop more advanced code generation and debugging methodologies. One direction is to extend MGDebugger to handle more complex bugs and code structures, such as multi-file projects and codebase with multiple dependencies. Another direction is to explore the collaboration of hierarchical code generation approaches such as Parsel (Zelikman et al., 2023) with hierarchical debugging, enabling end-to-end code generation and debugging systems. Furthermore, integrating MGDebugger into self-training systems to correct outputs from base models, then retraining the base models with the corrected data, could potentially improve their performance iteratively (Gulcehre et al., 2023).
Appendix A Detailed Hierarchical Decomposition Example
We provide the detailed illustration of the hierarchical decomposition process in MGDebugger, as shown in Figure 6, which has been simplified for illustration in Figure 1. For the original function “make_palindrome”, we decompose it into three minimal reusable subfunctions. And the relationships between subfunctions are naturally captured in the hierarchical structure based on the function calls. This hierarchical decomposition allows MGDebugger to systematically analyze and debug the code at different levels of granularity, leading to more effective identification and correction of bugs.
Figure 6: Detailed illustration of the hierarchical decomposition process in MGDebugger. The original code is decomposed into multiple sub-functions, each representing a significant step or logical block. The relationships between sub-functions are naturally captured in the hierarchical structure based on the function calls.
Appendix B HumanEvalFix Dataset
To access the ability of MGDebugger in debugging code with different types of bugs, we use the HumanEvalFix dataset (Muennighoff et al., 2023), which consists of 164 buggy functions across six programming languages, each provided with solutions and unit tests. For our experiments, we focus on the Python subset of the dataset. The buggy functions are categorized into six types of bugs: value misuse, missing logic, excess logic, operator misuse, variable misuse, and function misuse. Table 4 shows the distribution and explanations of these bug types within the HumanEvalFix dataset.
Table 4: Distribution and explanations of bugs in the HumanEvalFix dataset.
Bug Category Explanation Count
Value Misuse An incorrect value is used 44
Missing Logic Misses code needed to solve the problem 33
Excess Logic Contains excess code leading to mistakes 31
Operator Misuse An incorrect operator is used 25
Variable Misuse An incorrect variable is used 23
Function Misuse An incorrect function is used 8
Appendix C Implementation Details
We generate seed programs for HumanEval and MBPP using the BigCode Evaluation Harness framework [^3] https://github.com/bigcode-project/bigcode-evaluation-harness. The specific versions of models used in our experiments are DeepSeek-Coder-V2-Lite-Instruct [^4] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct, CodeQwen1.5-7B-Chat [^5] https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat, and Codestral-22B-v0.1666https://huggingface.co/TechxGenus/Codestral-22B-v0.1-GPTQ. All experiments are conducted on NVIDIA A100 GPUs with 80GB memory. During debugging, we use the vLLM engine [^7] https://github.com/vllm-project/vllm to serve the LLMs, setting the maximum token length according to each LLM’s max length. Following Zhong et al. (2024), we limit the maximum number of debugging iterations to 10 for all methods. Additionally, the sampling temperature is set to 0.8 in MGDebugger.
To obtain visible test cases for HumanEval and HumanEvalFix, we extract the given visible test cases from the task description. For MBPP, we use the first test case of each problem as the visible test case and use the rest as hidden test cases, in line with the settings referenced from Chen et al. (2023b) and Zhong et al. (2024).
Appendix D Debugging code with varying lengths on HumanEval and MBPP
To further demonstrate the robustness of MGDebugger in handling code of varying lengths, we present examples from MBPP and HumanEval. Similar to the examples provided for HumanEvalFix, we categorize the problems into short, medium, and long groups based on their code lengths, and we measure the repair success rates of MGDebugger and other baseline methods. All methods are built upon DeepSeek-Coder-V2-Lite. As is observed in Figure 8 and Figure 8, MGDebugger consistently outperforms other methods across different code lengths, especially in longer codes. This result demonstrates the scalability and robustness of MGDebugger in handling code of varying lengths and complexities again.
Figure 7: Repair success rate of different methods when debugging code of different lengths on HumanEval with DeepSeek-Coder. MGDebugger consistently performs the best across different code lengths.
Figure 8: Repair success rate of different methods when debugging code of different lengths in MBPP with DeepSeek-Coder. MGDebugger consistently performs the best across different code lengths.
Appendix E Impact of Debug Attempts on HumanEval and MBPP
We also investigate the impact of debug attempts on the cumulative repair success rate of MGDebugger and other methods on HumanEval and MBPP. As shown in Figure 10 and Figure 10, MGDebugger continues to improve with more debug attempts and achieves the highest success rate among all methods. Different from the results on HumanEvalFix that MGDebugger starts to ourperform other methods after the first attempt, MGDebugger significantly outperforms other methods from the beginning to the end on HumanEval and MBPP. This result highlights the effectiveness of MGDebugger in iterative and comprehensive debugging, making it a promising solution for complex and challenging code repair tasks.
Figure 9: Impact of debug attempts on the cumulative repair success rate of MGDebugger and other methods on HumanEval with DeepSeek-Coder. MGDebugger continues to improve with more debug attempts and outperforms other methods from the beginning to the end.
Figure 10: Impact of debug attempts on the cumulative repair success rate of MGDebugger and other methods on MBPP with DeepSeek-Coder. MGDebugger continues to improve with more debug attempts and outperforms other methods from the beginning to the end.
Appendix F Examples
We provide example code repairs for HumanEval, MBPP, and HumanEvalFix with DeepSeek-Coder-V2-Lite as the base model. The results of MGDebugger and baselines: Simple Feedback, Self-Edit, LDB (Block), LDB (Line), LDB (Function), Self-Debugging (Expl.), Self-Debugging (Trace) and Reflexion, are shown in the following tables. The buggy part in the original code is highlighted in yellow, and the repaired code is compared with the original buggy code, with changes highlighted in green if the repair passes the final test cases and in red if it fails. The functional comments in the solution code have been replaced with placeholders for brevity, as they are the same as those in the problem description.
The success of code repair often depends on initial solutions, as other methods typically change only a few lines of the original code, keeping the overall structure the same. This tendency to keep the structure of the initial solution may cause other methods to miss important parts of the code that are actually flawed. By contrast, by breaking down the code into smaller parts and looking at different levels of detail, our approach makes these seemingly correct but actually wrong parts easier to spot, as seen in the example of HumanEval. Also, the core of the code is often found in just a few lines, especially in math or dynamic programming problems. These sections usually need several steps of reasoning to understand and debug. By decomposing them into smaller subfunctions and simulating the execution to trace variable changes, MGDebugger can effectively identify and correct these bugs, as seen in the MBPP and HumanEvalFix examples.
F.1 Example on HumanEval
Info:  HumanEval/54
Problem:
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
$>$ same_chars('eabcdzzzz', 'dddzzzzzzzddeddabc')
True
$>$ same_chars('abcd', 'dddddddabc')
True
$>$ same_chars('dddddddabc', 'abcd')
True
$>$ same_chars('eabcd', 'dddddddabc')
False
$>$ same_chars('abcd', 'dddddddabce')
False
$>$ same_chars('eabcdzzzz', 'dddzzzzzzzddddabc')
False
Buggy Seed Solution:
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
return sorted(s0) == sorted(s1)
F.1.1 Simple Feedback
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
return Counter(s0) == Counter(s1)
Result:
F.1.2 Self-Edit
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
return Counter(s0) == Counter(s1)
Result:
F.1.3 LDB (Block)
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
if len(s0) != len(s1):
return False
for char in set(s0):
if s0.count(char) != s1.count(char):
return False
return True
Result:
F.1.4 LDB (Line)
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
return Counter(s0) == Counter(s1)
Result:
F.1.5 LDB (Function)
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
return Counter(s0) == Counter(s1)
Result:
F.1.6 Self-Debugging (Expl.)
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
return sorted(s0) == sorted(s1)
Result:
F.1.7 Self-Debugging (Trace)
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
return sorted(s0) == sorted(s1)
Result:
F.1.8 Reflexion
def same_chars(s0: str, s1: str):
Check if two words have the same characters.
if len(s0) > len(s1):
s0, s1 = s1, s0
it = iter(s1)
return all(char in it for char in s0)
Result:
F.1.9 MGDebugger
def same_chars(s0: str, s1: str) -> bool:
Check if two words have the same characters.
return sorted_chars_equal(s0, s1)
def sorted_chars_equal(s0: str, s1: str) -> bool:
"""Helper function to check if the sorted characters of two strings are equal."""
return sort_string(s0) == sort_string(s1)
def sort_string(s: str) -> list:
"""Helper function to sort the characters of a string."""
return sorted(set(s))
Result:
F.2 Example on MBPP
Info:  MBPP/185
Problem:
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
Buggy Seed Solution:
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = (1 - (b**2) / (4 * a * c))
return (x_focus, y_focus)
F.2.1 Simple Feedback
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = c - (b**2) / (4 * a)
return (x_focus, y_focus)
Result:
F.2.2 Self-Edit
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = c - (b**2) / (4 * a)
return (x_focus, y_focus)
Result:
F.2.3 LDB (Block)
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = (1 - (b**2) / (4 * a * c))
return (x_focus, y_focus)
Result:
F.2.4 LDB (Line)
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = c - (b**2) / (4 * a)
return (x_focus, y_focus)
Result:
F.2.5 LDB (Function)
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = (1 - (b**2) / (4 * a * c)) # Corrected formula for y_focus
return (x_focus, y_focus)
Result:
F.2.6 Self-Debugging (Expl.)
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = c - (b**2) / (4 * a)
return (x_focus, y_focus)
Result:
F.2.7 Self-Debugging (Trace)
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
# Calculate the discriminant
discriminant = b**2 - 4*a*c
# Calculate the x-coordinate of the focus
x_focus = -b / (2 * a)
# Calculate the y-coordinate of the focus
y_focus = (1 - discriminant) / (4 * a)
return (x_focus, y_focus)
Result:
F.2.8 Reflexion
def parabola_focus(a, b, c):
Write a function to find the focus of a parabola.
x_focus = -b / (2 * a)
y_focus = c - (b**2) / (4 * a)
return (x_focus, round(y_focus, 2)
Result:
F.2.9 MGDebugger
def parabola_focus(a, b, c):
"""Write a function to find the focus of a parabola."""
x_focus = calculate_x_focus(a, b)
y_focus = calculate_y_focus(a, b, c)
return (x_focus, y_focus)
def calculate_x_focus(a, b):
"""Calculate the x-coordinate of the focus of a parabola."""
return -b / (2 * a)
def calculate_y_focus(a, b, c):
"""Calculate the y-coordinate of the focus of a parabola."""
x_vertex = -b / (2 * a)
y_vertex = a * (x_vertex ** 2) + b * x_vertex + c
y_focus = y_vertex + 1 / (4 * a)
return y_focus
Result:
F.3 Example on HumanEvalFix
Info:  HumanEval/130
Problem:
def tri(n):
"""Everyone knows Fibonacci sequence, it was studied deeply by mathematicians in
the last couple centuries. However, what people don't know is Tribonacci sequence.
Tribonacci sequence is defined by the recurrence:
tri(1) = 3
tri(n) = 1 + n / 2, if n is even.
tri(n) = tri(n - 1) + tri(n - 2) + tri(n + 1), if n is odd.
For example:
tri(2) = 1 + (2 / 2) = 2
tri(4) = 3
tri(3) = tri(2) + tri(1) + tri(4)
= 2 + 3 + 3 = 8
You are given a non-negative integer number n, you have to a return a list of the
first n + 1 numbers of the Tribonacci sequence.
Examples:
tri(3) = [1, 3, 2, 8]
Buggy Seed Solution:
def tri(n):
if n == 0:
return [1]
my_tri = [1, 3]
for i in range(2, n + 1):
if i % 2 == 0:
my_tri.append(i / 2 + 1)
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + i + (i + 3) / 2)
return my_tri
F.3.1 Simple Feedback
def tri(n):
if n == 0:
return [1]
my_tri = [1, 3]
for i in range(2, n + 1):
if i % 2 == 0:
my_tri.append(1 + i / 2)
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + (i + 1) / 2)
return my_tri
Result:
F.3.2 Self-Edit
def tri(n):
if n == 0:
return [1]
if n == 1:
return [1, 3]
my_tri = [1, 3]
for i in range(2, n + 1):
if i % 2 == 0:
my_tri.append(1 + i // 2) # Use integer division
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + my_tri[i - 3])
return my_tri
Result:
F.3.3 LDB (Block)
def tri(n):
if n == 0:
return [1]
my_tri = [1, 3]
for i in range(2, n + 1):
if i % 2 == 0:
my_tri.append(1 + i / 2)
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + i + 1) / 2)
return my_tri
Result:
F.3.4 LDB (Line)
def tri(n):
if n == 0:
return [1]
my_tri = [1, 3, 2] # Initialize with the first three known values
for i in range(3, n + 1):
if i % 2 == 0:
my_tri.append(1 + i / 2)
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + my_tri[i - 3])
return my_tri
Result:
F.3.5 LDB (Function)
def tri(n):
if n == 0:
return [1]
my_tri = [1, 3, 2] # Initialize with the first three known values
for i in range(3, n + 1):
if i % 2 == 0:
my_tri.append(1 + i / 2)
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + my_tri[i - 3])
return my_tri
Result:
F.3.6 Self-Debugging (Expl.)
def tri(n):
if n == 0:
return [1]
if n == 1:
return [1, 3]
my_tri = [1, 3]
for i in range(2, n + 1):
if i % 2 == 0:
my_tri.append(1 + i // 2) # Use integer division
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + my_tri[i - 3])
return my_tri
Result:
F.3.7 Self-Debugging (Trace)
def tri(n):
if n == 0:
return [1]
my_tri = [1, 3, 2] # Initialize with tri(0), tri(1), and tri(2)
for i in range(3, n + 1):
if i % 2 == 0:
my_tri.append(1 + i // 2) # Use integer division
else:
my_tri.append(my_tri[i - 1] + my_tri[i - 2] + my_tri[i - 3])
return my_tri
Result:
F.3.8 Reflexion
def tri(n):
if n == 0:
return [1]
elif n == 1:
return [1, 3]
trib_sequence = [1, 3, 2])
for i in range(3, n + 1):
if i % 2 == 0:
trib_sequence.append(1 + i // 2)
else:
trib_sequence.append(trib_sequence[i - 1] + trib_sequence[i - 2] + trib_sequence[i - 3])
return trib_sequence
Result:
F.3.9 MGDebugger
def tri(n):
if n == 0:
return [1]
my_tri = [1, 3]
for i in range(2, n + 1):
if i % 2 == 0:
my_tri.append((1 + i / 2))
else:
my_tri.append(tri_odd(my_tri, i))
return my_tri
def tri_odd(my_tri, i):
"""Helper function to compute the Tribonacci sequence value for odd n."""
if i < 3:
return None # or return a default value
return my_tri[i - 1] + my_tri[i - 2] + tri_special(i)
def tri_special(n):
"""Helper function to compute the special value for odd n."""
return (n + 3) // 2
Result:
Appendix G Prompt Templates for MGDebugger
We provide prompt templates for the three main components of MGDebugger: hierarchical decomposition, test case generation, and sub-function debugging with LLM-simulated execution. These prompts are designed to guide the language model in generating the desired outputs for each step of the debugging process. They have been slightly modified for clarity and brevity, please refer to our codes if you need the exact prompt templates  [^8] https://github.com/YerbaPage/MGDebugger.
G.1 Prompt for Hierarchical Decomposition
Prompt Template for Hierarchical Decomposition SYSTEM_PROMPT: You are an AI assistant specialized in refactoring Python code into a tree-style hierarchical structure. USER_PROMPT: Convert the following Python code into a tree-style hierarchical structure with multiple levels of sub-functions. Each significant step or logical block should be its own function, and functions can call other sub-functions. Ensure that the main function calls these sub-functions in the correct order, creating a tree-like structure. Original Code: {code} Instruction: Please first analyze the codes step by step, and then provide the converted code in a Python code block. When providing the final converted code, make sure to include all the functions in a flattened format, where each function is defined separately.
G.2 Prompt for Test Case Generation
Prompt for Test Case Generation SYSTEM_PROMPT: You are an AI assistant specialized in analyzing Python functions and generating test cases. USER_PROMPT: Full Code: {full_code} Public Test Cases for the Main Function: {public_test_cases} Instruction: Please analyze how the {function_name} function is used within the main function and how it contributes to the expected outputs in the gold test cases. For each test case, you should analyze step-by-step based on both the input and the expected output of the main function, and then provide the corresponding input and expected output for the {function_name} function. Ensure that the generated test cases are consistent with the behavior expected in the public test cases.
G.3 Prompt for Debugging Subfunction
Prompt for Debugging Subfunction SYSTEM_PROMPT: You are an AI assistant helping to debug Python functions. USER_PROMPT: Debug the following Python function. The function is not passing all test cases. Analyze the code, identify the bug, and provide a fixed version of the function. Function Code: {function_code} Test Case Results: {test_case_results} Instruction: Please try to work as a Python interpreter to execute the code step-by-step. Identify the change of each variable as you ”run” the code line-by-line. Based on the execution trace, try to identify the bug and provide the final fixed code in a Python code block.

### v3

@@ 4.1. Case Study
We perform a qualitative analysis of how MGDebugger effectively identifies and corrects buggy parts compared to baseline methods. Figure 6 shows an example of debugging code snippets from the HumanEvalFix dataset using MGDebugger and other representative methods, with DeepSeek-Coder-V2-Lite as the backbone LLM. The original buggy solution computes the Collatz sequence with an incorrect computation logic of $n=n\times 2+1$. While other methods correct the computation to $n=n\times 3+1$, they introduce a new bug that misses the last “1” in the Collatz sequence. This is possibly because they get distracted by the need to filter odd numbers, and thus move the operation of appending the number to the results before updating $n$. MGDebugger excelled by decomposing the problem into distinct subfunctions: sequence generation and odd number filtering. By debugging each subfunction independently, MGDebugger ensured comprehensive error correction, including the subtle requirement of incorporating 1 into the Collatz sequence. This approach demonstrates MGDebugger’s ability to handle complex, multi-step problems more effectively than holistic debugging methods. Additionally, it highlights MGDebugger’s ability to not only fix bugs but also restructure code for enhanced clarity and correctness, demonstrating its potential in improving the quality of LLM-generated code.
4.2. Individual Component Analysis
To further validate the quality of individual components, we conducted manual verification on 38 examples from HumanEval, including 29 successfully debugged and 9 failed cases, using DeepSeek-Coder-V2-Lite. Our analysis reveals three key findings: 1) Code Decomposition: All 38 decomposed versions are functionally equivalent to the original functions, demonstrating 100% semantic correctness in our decomposition process. Since the overall codes have been provided, it’s less likely that the model includes new bugs when transforming the code into subfunctions, demonstrating the high quality of our LLM-based decomposition approach. 2) Test Case Generation: 291 out of 312 generated test cases, representing 93.2% accuracy, were correct when the model is instructed to strictly follow the public test cases when generating test cases for each subfunction. Since our approach derives test cases exclusively from the original public test cases, with a mean test coverage of 71.3%, mutation testing scores of 9.2% and test diversity of 28.6%, we maintain the inherent quality characteristics of the original test suite while enhancing debugging effectiveness. Additionally, in a supplementary experiment on 9 failed debugging cases, providing additional test cases that exposed errors enabled successful debugging of 2 additional functions, indicating potential for further improvement with enhanced test diversity and coverage. 3) Simulated Execution: The LLM correctly predicted 389 of 413 breakpoints, achieving 94.2% accuracy, with failures occurring only in cases requiring complex multi-step reasoning. The high accuracy aligns well with existing studies (Zheng et al., 2024; Ni et al., 2024). Notably, when we manually corrected the prediction errors, 2 of the 9 failed debugging cases could be successfully fixed, indicating potential for further improvement with enhanced simulation accuracy. According to our manual analysis, the high quality of each component in MGDebugger contributes to the overall effectiveness, as demonstrated by the ablation results. This suggests our approach is robust and that further improvements in component quality could lead to even better results.
4.3. Bug Granularity Analysis
To validate our main claim that hierarchical debugging helps find bugs at different granularity levels, we manually analyzed the 23 problems successfully fixed by MGDebugger using DeepSeek-Coder on HumanEval. Our analysis reveals that 13 out of 23 cases, representing 56.5%, involved low-level bugs such as syntactic errors, API misuse, incorrect operators, and variable name errors, while 10 out of 23 cases, representing 43.5%, involved high-level bugs including algorithmic logic errors, incorrect problem understanding, and missing core logic. This distribution demonstrates that MGDebugger effectively addresses both syntactic and semantic faults across different granularity levels, validating our hierarchical debugging strategy for handling diverse error types. The ability to handle both categories showcases MGDebugger’s versatility across various debugging scenarios.
4.4. Evaluation on Unseen Data
Since HumanEval and Defects4J are potentially contained in training data of LLMs, potential data leakage could affect our evaluation. To mitigate this threat, we have employed several strategies: (1) we utilized the HumanEvalFix dataset, which consists of human-crafted bugs created after the training cutoff dates of Codestral; (2) our experimental design inherently reduces the impact of potential data leakage by focusing exclusively on test cases where the No-Debugging LLM fails to produce correct solutions (Olausson et al., 2023; Chen et al., 2023b). If data leakage had a strong effect for a particular case, the model would likely have generated a correct solution on its first attempt, and such cases would not be included in our debugging evaluation set.
To further address potential data leakage concerns, we conducted additional experiments on LiveCodeBench (Jain et al., 2025), which contains programming problems released after LLM training cutoffs. Using 131 problems from the February-May 2025 evaluation window with GPT-4o, MGDebugger achieves 36.6% accuracy with 10.8% repair success rate, outperforming all baselines by +7.6% over No-Debugging and +3.0% over Reflexion (Table 6). These results also confirm that MGDebugger’s improvements are not artifacts of data leakage but represent genuine debugging capabilities on completely unseen data.
Table 6. Results on LiveCodeBench with GPT-4o.
Method Acc. (%) RSR (%)
No-Debugging 29.0
Self-Debugging (Expl.) 32.8 (+3.8) 4.3
Reflexion 33.6 (+4.6) 5.4
MGDebugger 36.6 (+7.6) 10.8
4.5. Threats to Validity
While our evaluation is comprehensive, several threats to validity should be acknowledged. 1) The first is the risk of limited generalizability across programming languages and domains. To mitigate this concern, we have evaluated MGDebugger across four distinct benchmarks spanning both Python and Java languages. This comprehensive validation across different language contexts strengthens our confidence in the approach’s broad applicability. 2) Our method’s performance may vary with different LLM architectures. To address this concern, we conducted experiments with various open-source LLMs of different sizes as well as several commercial models. The consistent improvements observed across this diverse set of models suggest that MGDebugger’s benefits are robust across different model architectures and capabilities. 3) For our evaluation on Defects4J, an internal threat arises from the manual validation process used to determine the correctness of the generated patches. To mitigate this concern, we conducted thorough examinations of each patch, following established validation practices from prior work in automated program repair (Xia and Zhang, 2024; Kong et al., 2025; Lin et al., 2024; Yin et al., 2024; Zhang et al., 2024a).
Despite these potential limitations, our extensive evaluation across multiple dimensions—including various models, datasets, bug types, and code complexities—provides strong evidence that MGDebugger represents a significant advancement in debugging LLM-generated codes. The consistency of our results and the magnitude of improvement over existing methods indicate broad applicability and robust performance that is likely to extend to many real-world debugging scenarios.
5. Related Work
5.1. Large Language Models for Code
Recent models such as GPT4 (OpenAI, 2023), Codestral (Mistral AI team, 2024), and DeepSeek-Coder (AI, 2024) have advanced code generation through instruction tuning and reinforcement learning from human feedback (RLHF) with mixed code and natural language data (Ziegler et al., 2020; Rafailov et al., 2023; Anthropic, 2024). Code generation with LLMs has been enhanced by various techniques. Some approaches focus on improving the quality of generated code using planning algorithms, transitioning from outlines to detailed implementations (Zhang et al., 2022; Yao et al., 2023; Zelikman et al., 2023; Zheng et al., 2023). Other methods sample multiple programs from the same LLM and rank them to identify the best one (Chen et al., 2023a, 2022; Ni et al., 2023). Additionally, some works leverage multi-agent collaboration frameworks to enhance code generation quality (Zhang et al., 2024b; Huang et al., 2023a; Dong et al., 2024). These approaches aim to optimize the production of correct code from the outset. By contrast, MGDebugger targets the post-generation phase, focusing on debugging and fixing errors that inevitably arise during the code generation process.
5.2. Repairing LLM-Generated Code
Program repair is a critical aspect of software development, aiming to automatically identify and fix bugs in code (Just et al., 2014; Gupta et al., 2020; Yasunaga and Liang, 2021; Shao et al., 2025). There are two main streams of research in repairing code generated by LLMs: (1) training models to repair code (Huang et al., 2023b; Jiang et al., 2024; Ding et al., 2024; Zheng et al., 2024; Moon et al., 2024; Kumar et al., 2024) and (2) providing external feedback to the raw pretrained models to fix code (Jiang et al., 2023b; Chen et al., 2023b; Olausson et al., 2023; Zhong et al., 2024; Hu et al., 2024). By contrast to previous work that trains separate models for code repair (Ding et al., 2024; Zheng et al., 2024; Moon et al., 2024), MGDebugger does not require task-specific retraining but takes advantage of the inherent capabilities of pretrained LLMs. This flexibility allows MGDebugger to operate in zero-shot settings, offering a lightweight and scalable alternative. Furthermore, exploring the ability of LLMs to fix their own code is a promising direction for self-improvement training of the LLMs (Burns et al., 2023).
MGDebugger falls under the category of work that leverages pretrained models to fix code by reasoning with external feedback. Several recent methods (Zhang et al., 2023; Olausson et al., 2023; Bouzenia et al., 2024; Lee et al., 2024) utilize execution results from test cases to guide LLMs in code correction. More recent works have explored advanced debugging techniques utilizing LLM’s reasoning ability. Reflexion (Shinn et al., 2023) prompts LLMs to reflect on the generated code and uses a memory buffer for iterative refinement. Self-Debugging (Chen et al., 2023b) prompts LLMs to explain or dry run generated programs, known as rubber duck debugging. LDB (Zhong et al., 2024) segments programs into basic blocks, tracking variable values during runtime after each block to verify the correctness against the task description. Although these methods incorporate detailed execution feedback and iterative refinement, they treat the whole function as a single unit and perform sequential debugging, limiting their effectiveness with complex code (Chen et al., 2023b; Hossain et al., 2024). MGDebugger addresses this issue by introducing a hierarchical approach, debugging from low-level errors to high-level flaws. This method ensures a more systematic and accurate debugging process, especially for complex and multifunctional systems.
5.3. Automated Program Repair
Traditional APR techniques can be broadly classified into three categories: template-based (Liu et al., 2019; Zhang et al., 2024a), heuristic-based (Le Goues et al., 2012; Kim et al., 2024) and constraint based (Le et al., 2017; Mechtaev et al., 2016; Zhu et al., 2023). While these approaches have demonstrated effectiveness, they often face challenges in scalability and patch diversity. The advent of deep learning has shifted the focus to learning-based approaches, particularly those framing program repair as a neural machine translation task (Lin et al., 2024; Li et al., 2020; Lutellier et al., 2020; Jiang et al., 2023a). These methods use encoder-decoder architectures to translate buggy code into correct versions, achieving notable success but remaining limited by the quality and quantity of training data (Yin et al., 2024). More recently, LLM-based APR tools, powered by models like ChatGPT and Codex, have achieved superior fix rates on standard benchmarks like Defects4J (Just et al., 2014) by generating diverse and contextually relevant patches (Xia et al., 2023; Xia and Zhang, 2024; Kong et al., 2025; Yin et al., 2024; Hossain et al., 2024; Xin et al., 2024; Zhang et al., 2025).
MGDebugger differs from existing APR approaches in both objective and methodology. MGDebugger focuses on enabling LLMs to correct their own mistakes from code generation, working towards more reliable AI code generation systems. Methodologically, MGDebugger introduces a novel bottom-up hierarchical debugging strategy combined with test case generation and LLM-simulated execution, unleashing the potential of LLMs’ reasoning capabilities.
6. Conclusion
@@
Future work can build upon this foundation to develop more advanced code generation and debugging methodologies. One direction is to extend MGDebugger to handle more complex bugs and code structures, such as multi-file projects and codebase with multiple dependencies. Another direction is to explore the collaboration of hierarchical code generation approaches with hierarchical debugging, enabling end-to-end code generation and debugging systems. Furthermore, integrating MGDebugger into self-training systems to correct outputs from base models, then retraining the base models with the corrected data, could potentially improve their performance iteratively.